OpenStudy (anonymous):

Question 2.2. How many solutions are there to the following system of equations? 2x + 5y = –11 10x + 25y = –55 (Points : 4) 0 1 2 infinitely many

OpenStudy (texaschic101):

try dividing the 2nd equation by 5 and what do you get ?

OpenStudy (texaschic101):

10x/5 + 25y/5 = -55/5 = 2x + 5y = -11 same equation as first one...therefore, they are the same line making them have infinite solutions